OpenStudy (anonymous):

Solve the following problem for the roots by using the quadratic formula. 2(6 - x) = x(x + 5)

OpenStudy (anonymous):

2(6 - x) = x(x + 5) i.e. 12 - 2x = x^2 +5 i.e. x^2 +5 +2x -12=0 i.e. x^2 +2x -7=0 comparing it with general eq of quadratic eq i.e ax^2 +bx +c = 0, we get a= 1, b= 2, c= -7 Quadratic eq formula is given as: x=(-b±√(b^2-4ac))/2a x=(-2±√{2^2-4*1*(-7)}/2*1 x=(-2±√{4+28}/2 x=(-2±√{32}/2 x=(-2±4√{2}/2 x=(-1±2√{2} which are the required roots of x.